In a nuclear reactor, each atom of uranium (of atomic mass 235 u) releases about 200 MeV when it fissions. What is the change in mass when 1.30 kg of uranium-235 is fissioned?

Answer in Kg
Answer #1

Solution :-

for each atom of Urnaium

energy emitted = 200 MeV

for 1.30 kg of U-235

energy emitted = number of atoms * energy per atom

energy emitted = 200 * (1.30×10^3/235) * 6.022 *10^23 MeV

let the change in mass is m

as emergy emitted = m * c^2

200 * (1.30×10^3/235) * 6.022 *10^23 * 1.602 *10^-19 * 10^6 = m * (3 *10^8)^2

solving for m

m = 0.001185 Kg

the change in mass is 0.001185 kg